OpenStudy (2yoututu11):

What is the volume of the cube? A. 6 cubic inches B. 8 cubic inches C. 12 cubic inches D. 16 cubic inches

OpenStudy (anonymous):

considering that the volumes formula is V = a^3 since a = 2 V = 2^3 V = 8 the answer would be B; 8 cubic inches